OpenStudy (anonymous):

x g(x) -4 -4 -2 -2 2 2 @sarahusher

OpenStudy (anonymous):

again x=g(x) so i presume you have something different that it might intersect? :)

OpenStudy (anonymous):

How can one halfx − 5 = one thirdx + 6 be set up as a system of equations? sorry just ignore the top

OpenStudy (anonymous):

oh okay haha so we have \[\frac{ 1 }{ 2 }x-5=\frac{ 1 }{ 3 }x + 6\] okay so if they equal each other, we know they equal some value, let's call it y so we know y=(x/2)-5 and y=(x/3)+6 right?

OpenStudy (anonymous):

ok so do we put it beside each other?

OpenStudy (anonymous):

you don't have two the two equations listed as they are is enough that they satisfy a 'system of equations' as the questions asks

OpenStudy (anonymous):

right im soo sorry

OpenStudy (anonymous):

don't apologise dear! :)

OpenStudy (anonymous):

2y + x = −10 3y + x = 18 2y + 2x = −10 3y + 3x = 18 2y − x = −10 3y − x = 18 2y − 2x = −10 3y − 3x = 18

OpenStudy (anonymous):

which one would it be then @sarahusher

OpenStudy (anonymous):

so if we have:\[y=\frac{ x }{ 2 }-5\] and \[y=\frac{ x }{ 3 }+6\] you can multiply the top one by two (so everything is whole numbers) and the bottom one by 3, what do you get then?

OpenStudy (anonymous):

2-10 3+18

OpenStudy (anonymous):

is that right @sarahusher

OpenStudy (anonymous):

right idea but you have to include the 'y' component so it would be 2y=x-10 3y=x+18

OpenStudy (anonymous):

right i got you thanks

OpenStudy (anonymous):

you're welcome